A 800 N mass is hanging on a 10 m length board. The board weighs 200 N and is supported by a wall and a rope. The rope and the board has an angle between them of 30º and the rope is 2 m from the right end of the board. FInd the tension in the rope and the force by the wall on a board.


1
Expert's answer
2021-03-17T16:54:41-0400


Equilibrium of moments relative to point O:


"-10W_2-5W_1+8T\text{ sin}30°=0,\\
T=2250\text{ N}."

Equilibrium of forces:


"y: T\text{ sin}30°+R_y-W_1-W_2=0,\\
x: R_x-T\text{ cos}30°=0.\\\space\\
R_y=-125\text{ N},\\
R_x=1950\text{ N},\\
R=\sqrt{R_x^2+R_y^2}=1953\text{ N}
."